How much time is needed for a sample of Pd-100 to lose 93.75% of its original amount? Pd-100 has a half-life of 3.634 days.Give an exponential model as a function t.
Answer by jorel1380(3719)

Let n be the original amount of pd-100. If it loses 93.75% of its' original amount, then the amount left equals 100-93.75=6.25%. So:
.0625=(.5)^t/3.634
.5^4=.5^t/3.634
t/3.634=4
t=14.536 or
log .0625=log .5^t/3.634
log .0625=t/3.634*log .5
(log .0625/log .5)*3.634=t
t=14.536 days. ☺☺☺☺